### v1

# Some Bayesian Perspectives on Clinical Trials

Abstract
We provide a Bayesian perspective on three interconnected aspects of clinical trial design: prior specification, sequential adaptive allocation, and decision-theoretic optimization. For prior specification, we argue that treatment effects in clinical trials are known a priori to be small, rendering default noninformative priors such as Jeffreys’ prior inappropriate; priors calibrated to historical effect sizes or LD50 relationships are both more honest and more efficient. For sequential design, we show how Thompson’s (1933) probability-matching rule connects to modern adaptive randomization, and how backward induction on sufficient statistics—following Christen and Nakamura [2003] and Carlin et al. [1998]—reduces the seemingly intractable infinite-horizon stopping problem to a finite table. For trial optimization, we review the utility-based framework of Thall and Cook [2004] that jointly models efficacy and toxicity, enabling dose-finding designs that maximize patient benefit rather than merely controlling error rates. We illustrate these ideas through the ECMO trial, the CALGB 49907 breast cancer trial, and modern platform trials, and discuss the 2026 FDA draft guidance on Bayesian methodology.
1 Introduction
The standard paradigm for clinical trial design—fix the sample size, randomize equally, analyze once at the end—was built to control the probability of approving an ineffective drug. It succeeded at that goal. But the framework treats patients enrolled during the trial as expendable inputs to a statistical procedure whose conclusions matter only afterward. If early data show one treatment to be clearly inferior, patients continue receiving it. If the trial is underpowered, patients’ participation generates inconclusive evidence.
Bayesian methods offer a different organizing principle: probability is a measure of uncertainty, updated coherently as data arrive. This principle has three consequences for clinical trials that we develop in this paper:
  1. 1.
    Priors encode what is already known. Treatment effects in clinical trials are small—typically a few percentage points of absolute risk reduction. This prior knowledge, ignored by noninformative analyses, can improve both efficiency and calibration when formalized correctly.
  2. 2.
    Sequential decisions can be made optimally. When utility depends on sufficient statistics, the infinite-horizon stopping problem reduces to a computable table via backward induction. Combined with Thompson’s (1933) probability-matching rule, this yields adaptive designs that balance learning against treating.
  3. 3.
    Trial design is a decision problem. The choice of dose, the rule for stopping, and the allocation of patients are decisions under uncertainty. A utility function that jointly values efficacy, toxicity, and information gain provides a coherent framework for making them.
We are not the first to advocate Bayesian clinical trials. Novick and Grizzle [1965] introduced Bayesian analysis of clinical data using conjugate priors for categorical outcomes. Berry et al. [2010] provide a comprehensive treatment of Bayesian adaptive methods. Berry [2025] reviews the state of the art. Our contribution is to connect three specific arguments—that default priors are wrong, that backward induction is tractable, and that utility-based design unifies ad hoc adaptations—and to illustrate them through trials where these ideas made a material difference.
Section 2 discusses prior specification. Section 3 covers sequential adaptive designs. Section 4 develops the decision-theoretic framework. Section 5 presents case studies. Section 6 discusses the regulatory landscape. Section 7 concludes.
2 The Role of Priors
The prior distribution is the defining feature of Bayesian inference and the source of most controversy surrounding it. In clinical trials, the controversy is misplaced: we know far more about treatment effects before a trial begins than a noninformative prior would suggest.
2.1 Why default priors fail for treatment effects
Consider a two-arm trial comparing a new treatment to control, with binary outcome (response/no response). Let $\theta_{1}$ and $\theta_{0}$ denote the response probabilities under treatment and control, respectively. The treatment effect is $\delta=\theta_{1}-\theta_{0}$.
Jeffreys’ prior for the binomial proportion is $\text{Beta}(1/2,1/2)$, which places substantial mass near 0 and 1. Under independent Jeffreys priors for $\theta_{1}$ and $\theta_{0}$, the implied prior on $\delta$ is U-shaped, concentrating mass near $\pm 1$. This contradicts everything known about drug development: treatment effects are small. Among drugs that reach Phase III, the median absolute risk reduction is on the order of 5–10 percentage points [Berry, 2012]. A prior that places most mass on treatment effects near $\pm 100\%$ is not noninformative—it is misinformative.
The uniform prior $\text{Beta}(1,1)$ is less extreme but still problematic. It implies that a response rate of 95% is as likely as 50% a priori, and that the treatment effect $\delta$ has prior standard deviation of approximately 0.41—far larger than any plausible effect in most therapeutic areas.
Figure 1 makes the problem visible: the implied prior on $\delta$ under Jeffreys priors is U-shaped, placing most mass on extreme effects near $\pm 1$; under uniform priors it is triangular, still far too diffuse; only an informative prior concentrates mass where clinical effects actually live.
Figure 1: Implied prior density on the treatment effect $\delta=\theta_{1}-\theta_{0}$ under three choices of independent priors for the success probabilities $\theta_{0}$ and $\theta_{1}$. The shaded band marks the range of typical clinical effects ($|\delta|\leq 0.15$). Jeffreys priors (red) concentrate mass near $\pm 1$; uniform priors (blue) spread mass broadly; an informative prior (grey) concentrates mass where treatment effects are known to lie.
Novick and Grizzle [1965] recognized this issue and proposed “logical probability” priors calibrated to prior data. Their natural conjugate Bayes densities express prior information as equivalent to $(m,z)$ prior observations with $z$ successes, so the prior $\text{Beta}(z+1,m-z+1)$ has the interpretation of $m$ prior patients. This parameterization makes the prior’s influence transparent: it adds $m$ pseudo-observations to the data. The effective sample size concept was later formalized by Morita et al. [2008].
2.2 Informative priors from historical data
A more principled approach uses historical data to construct informative priors. Racine et al. [1986] demonstrated this systematically across four areas of pharmaceutical research at Ciba-Geigy: LD50 experiments, crossover trials, bioequivalence assessment, and population pharmacokinetics. For LD50 experiments, where small sample sizes cause fiducial intervals to degenerate to the entire real line, Bayesian priors calibrated to toxicity class boundaries produced posterior densities that were both well-defined and practically useful. For population pharmacokinetic models—hierarchical nonlinear random effects models where individual parameters $\theta_{i}$ vary across patients—the Bayesian framework borrows strength across individuals, enabling prediction of a new patient’s drug concentration profile from as few as two blood draws when combined with population data. If a drug’s mechanism of action is understood, prior knowledge about the control arm response rate is often substantial. For dose-response relationships, the LD50 (the dose producing 50% response) provides a natural anchor: a prior on the dose-response curve centered at the LD50 with variance reflecting uncertainty about the curve’s slope encodes genuine pharmacological knowledge.
A fundamental difficulty in specifying priors for dose-response models is that regression coefficients are abstract and model-dependent: the meaning of $\beta$ in a logistic model differs from its meaning in a probit model, so eliciting priors directly on $\beta$ is unnatural. Bedrick et al. [1996] resolved this by introducing conditional means priors: the investigator specifies prior distributions on observable quantities—the probability of response at selected covariate values—and the prior on $\beta$ is induced through the link function. For a binomial regression with $p$ predictors, one specifies independent $\text{Beta}(a_{1i},a_{2i})$ priors on the success probabilities $\tilde{m}_{i}$ at $p$ chosen covariate vectors $\tilde{\mathbf{x}}_{i}$, and the induced prior on $\beta$ takes the form
$\pi(\beta)\propto\prod_{i=1}^{p}F(\tilde{\mathbf{x}}_{i}^{\prime}\beta)^{a_{1i}-1}[1-F(\tilde{\mathbf{x}}_{i}^{\prime}\beta)]^{a_{2i}-1}f(\tilde{\mathbf{x}}_{i}^{\prime}\beta),$ (1)
where $F$ is the CDF associated with the link function and $f$ its density. The approach is concrete. In the Challenger O-ring analysis, Bedrick et al. [1996] modeled the probability that any O-ring fails as a function of temperature via $g(m_{i})=\beta_{0}+\beta_{1}(\tau_{i}-\bar{\tau})$ and specified independent beta priors on the failure probabilities at two temperatures: $\tilde{m}_{1}\sim\text{Beta}(1,0.577)$ at $55^{\circ}$F (giving $\Pr[\tilde{m}_{1}>1/2]=2/3$, reflecting a belief that failure is likely at low temperatures) and $\tilde{m}_{2}\sim\text{Beta}(0.577,1)$ at $75^{\circ}$F (giving $\Pr[\tilde{m}_{2}<1/2]=2/3$). These priors on observable failure rates are easy to elicit; the induced prior on $(\beta_{0},\beta_{1})$ via (1) then depends on the link function, so the same scientific beliefs produce different coefficient priors for logistic versus probit models without requiring the investigator to think about model-dependent abstractions. The idea traces back to Racine et al. [1986], who used independent beta priors on two success probabilities to induce bivariate priors on probit regression coefficients in bioequivalence studies at Ciba-Geigy.
For clinical trials, the translation is direct: a clinician can state beliefs about response rates at clinically meaningful doses (“at 10 mg, I expect 30–50% response; at 50 mg, perhaps 60–80%”) and the conditional means prior machinery converts these into a coherent prior on regression coefficients for whatever dose-response model is chosen. Figure 2 illustrates this for a dose-response trial: independent Beta priors on response probabilities at two doses (left panel) induce through the logistic link a proper, concentrated joint prior on the intercept and slope (right panel).
Figure 2: Conditional means prior for a dose-response model. Left: the clinician specifies Beta priors on the response probability at a low dose (10 mg, blue) and a high dose (50 mg, red). Right: the induced joint prior on the logistic regression coefficients $(\beta_{0},\beta_{1})$ via (1). The prior on coefficients is proper and concentrated, even though the clinician never reasoned about regression parameters directly.
For logistic regression specifically, the induced prior is equivalent to a data augmentation prior—pseudo-observations at the design points with fractional sample sizes—connecting directly to the Pólya-Gamma scheme below. Hanson et al. [2014] extended this through informative $g$-priors for logistic regression, where the investigator specifies only a prior on the overall population success probability and the $g$-prior structure induces a proper joint prior on all coefficients. Together with Chaloner [1993], who characterized optimal Bayesian designs for dose-response experiments under such priors, these methods give the trialist a complete toolkit: specify beliefs about response rates at clinically meaningful doses, let the link function induce priors on model parameters, and design the experiment to maximize information about the LD50 or other quantities of interest.
The FDA’s 2026 draft guidance on Bayesian methodology [Food and Drug Administration, 2026] identifies several settings where informative priors from historical data have supported regulatory submissions: borrowing from previous clinical trials of the same drug, augmenting concurrent controls with external data, pediatric extrapolation from adult trials, and borrowing across similar disease subtypes. In each case, the prior’s influence must be quantified—for example, through its effective sample size—and sensitivity analyses must demonstrate robustness to prior-data conflict.
Dynamic borrowing methods, where the degree of information transfer depends on observed consistency between historical and current data, mitigate the risk of static priors that do not respond to conflict. The Bayesian hierarchical framework provides a natural mechanism: the between-study variance parameter $\tau^{2}$ governs borrowing, with large $\tau^{2}$ producing skeptical priors that are quickly overwhelmed by current data.
2.3 Computational advances: Pólya-Gamma augmentation
Many clinical trial endpoints are binary or categorical, leading to logistic regression models for treatment effects adjusted for covariates. Bayesian inference for logistic regression historically required Metropolis-Hastings sampling, which is slow and requires careful tuning.
Polson et al. [2013] introduced the Pólya-Gamma data augmentation scheme, which renders the posterior for logistic regression conditionally Gaussian. The key identity is:
$\frac{(e^{\psi})^{a}}{(1+e^{\psi})^{b}}=2^{-b}e^{\kappa\psi}\int_{0}^{\infty}e^{-\omega\psi^{2}/2}p(\omega)\,d\omega,$ (2)
where $\kappa=a-b/2$ and $\omega\sim\text{PG}(b,0)$ is a Pólya-Gamma random variable. Conditional on $\omega$, the log-odds $\psi$ has a Gaussian full conditional, enabling a simple Gibbs sampler that alternates between drawing $\omega$ from a Pólya-Gamma distribution and drawing regression coefficients from a multivariate Gaussian.
For clinical trials with binary endpoints and moderate covariate adjustment, this scheme produces exact posterior samples with no tuning parameters, making real-time Bayesian monitoring computationally straightforward. The approach extends naturally to multinomial and negative binomial models.
3 Sequential Adaptive Designs
The ethical case for adaptive designs is simple: if accumulating data show one treatment to be inferior, continuing to assign patients to it is unjustifiable. The statistical case is that sequential allocation can improve efficiency, reducing the number of patients needed to reach a conclusion.
The frequentist approach to interim monitoring is group sequential testing. Pocock [1977] proposed equal critical values at each interim look; O’Brien and Fleming [1979] proposed conservative early boundaries that spend little alpha initially. Lan and DeMets [1983] unified these through the alpha-spending function $\alpha^{*}(t)$, which allocates the overall type-I error rate across information fractions $t\in[0,1]$. The spending function need not be specified until the interim analysis is conducted, giving the trialist flexibility in the timing and number of looks.
The Bayesian sequential framework differs in three respects. First, it conditions on the data actually observed rather than on a pre-specified schedule of looks, so inference is valid regardless of when or why the data were examined. Second, the stopping rule is based on predictive or posterior probabilities that answer the clinical question directly (“what is the probability that this treatment is superior?”) rather than on $p$-values adjusted for multiplicity. Third, the Bayesian framework extends naturally to response-adaptive randomization and multi-arm designs, whereas group sequential methods are largely restricted to fixed-allocation two-arm trials with pre-planned interim analyses.
3.1 Thompson sampling
Thompson [1933] proposed the earliest response-adaptive randomization rule. Given two treatments with unknown success probabilities $\tilde{p}_{1}$ and $\tilde{p}_{2}$, and prior data consisting of $r_{i}$ successes in $n_{i}$ patients for treatment $i$, Thompson computed
$P(\tilde{p}_{2}>\tilde{p}_{1}\mid\text{data})=\frac{\sum_{\alpha=0}^{r_{2}}\binom{r_{1}+r_{2}-\alpha}{r_{1}}\binom{s_{1}+s_{2}+1+\alpha}{s_{1}}}{\binom{n_{1}+n_{2}+2}{n_{1}+1}}$, (3)
where $s_{i}=n_{i}-r_{i}$, and proposed assigning the next patient to treatment 2 with this probability. This “probability matching” rule assigns patients to the treatment that is more likely to be better, in proportion to the evidence.
Thompson’s motivation was explicitly ethical: “a considerable saving of individuals otherwise sacrificed to the inferior treatment might be effected.” Under uniform priors, the posterior for each treatment’s success rate is $\text{Beta}(r_{i}+1,s_{i}+1)$, and the allocation probabilities in (3) follow from the integral $\int_{0}^{1}\int_{0}^{\theta_{2}}f(\theta_{2}|\text{data})f(\theta_{1}|\text{data})\,d\theta_{1}\,d\theta_{2}$.
Thompson sampling was rediscovered by the technology industry in the 2000s and is now the dominant approach to online optimization problems—website design, advertisement placement, recommendation systems, and cloud service configuration [Scott, 2010]. Advances in Bayesian computation made randomized probability matching easy to implement for virtually any payoff distribution. If Thompson sampling has become the standard for allocating web traffic to competing designs, where the stakes are advertising revenue, it should a fortiori be the standard for allocating patients to competing treatments, where the stakes are human lives.
Figure 3 shows simulated Thompson sampling paths for a two-arm trial with control success rate $p_{0}=0.30$ and treatment success rate $p_{1}=0.45$. The allocation proportion drifts above the equal-randomization line as evidence accumulates, assigning more patients to the superior arm.
Figure 3: Thompson sampling in a two-arm trial ($p_{0}=0.30$, $p_{1}=0.45$). Each curve shows the running proportion of patients assigned to the treatment arm across eight independent simulations. Under equal randomization (dashed line), half of all patients receive the inferior control. Thompson sampling progressively shifts allocation toward the superior arm as evidence accumulates.
The Thall-Wathen generalization raises the posterior probabilities to a power $\tau$ before normalizing:
$r_{k,n}=\frac{[p_{k,n}]^{\tau}}{\sum_{j=1}^{K}[p_{j,n}]^{\tau}},$ (4)
where $p_{k,n}=\Pr(\pi_{k}>\pi_{j}\text{ for all }j\neq k\mid D_{n})$ is the posterior probability that arm $k$ is best after $n$ patients, and $K$ is the number of arms. When $\tau=0$, this reduces to equal randomization; when $\tau\to\infty$, it assigns all patients to the current leader. The tuning parameter $\tau$ controls the exploration-exploitation tradeoff inherent in the multi-armed bandit formulation of clinical trials.
3.2 Predictive probabilities and stopping rules
A trial’s most consequential decision is when to stop. Bayesian predictive probabilities answer the question trial teams actually face: given current data, what is the probability that the trial will reach a definitive conclusion if continued? The idea has an early precedent in bioequivalence testing: Racine-Poon et al. [1987] proposed a two-stage procedure in which first-stage data are used to compute the predictive probability of establishing bioequivalence in a second stage, and the second-stage sample size $n_{2}$ is chosen—or the study abandoned—accordingly. Their simulations showed that the two-stage design roughly doubles the success rate of single-stage studies at a comparable average sample size.
Let $Z_{n}$ denote the test statistic at the interim analysis with information fraction $r=I_{n}/I_{N}$, where $I_{n}$ is the current information and $I_{N}$ is the information at the planned final analysis. Marion et al. [2025] show that the predictive probability of trial success can be approximated as
$\text{PP}(p_{n},r,\alpha)=\Phi\left(\frac{\Phi^{-1}(1-p_{n})-\Phi^{-1}(1-\alpha)\sqrt{r}}{\sqrt{1-r}}\right),$ (5)
where $p_{n}$ is the current $p$-value (or $1-\mathcal{P}_{n}$ for Bayesian posterior probability $\mathcal{P}_{n}$) and $\alpha$ is the significance threshold. This approximation is exact when the test statistic is Gaussian with known information fraction and a flat prior, and accurate across a range of endpoint types—continuous, binary, time-to-event, and ordinal—as verified by simulation against full Monte Carlo imputation [Marion et al., 2025].
The computational advantage is substantial: where full predictive probability calculations require nested simulations (simulating future data within each simulated interim analysis), the approximation in (5) requires only the current $p$-value and the information fraction. This enables trial simulation with orders-of-magnitude speedup.
3.3 Backward induction on sufficient statistics
The optimal sequential stopping rule can in principle be found by backward induction: starting at the maximum horizon, compute the optimal decision for every possible data configuration, then work backward. For general data, the decision tree grows exponentially and the computation is intractable.
Christen and Nakamura [2003], building on Berry et al. [2000] and Carlin et al. [1998], identified a key structural insight: when the utility function depends on data only through a sufficient statistic, the exponentially growing decision tree collapses to a polynomial-size table.
Consider a sequential experiment with horizon $T$, data $\mathbf{X}_{t}=(X_{1},\ldots,X_{n+t})$ after $t$ additional observations, and utility for stopping at time $t$:
$u(1,\mathbf{X}_{t})=a\,S(t)+b(T-t),$ (6)
where $S(t)=\sum_{l=1}^{n+t}X_{l}$ is the running total and $a,b$ are reward and cost parameters. The utility of continuing is the expected optimal utility at the next step:
$u(0,\mathbf{x}_{t})=\sum_{\mathbf{x}_{t+1}}U(\mathbf{x}_{t+1})\,P(\mathbf{X}_{t+1}=\mathbf{x}_{t+1}\mid\mathbf{X}_{t}=\mathbf{x}_{t}),$ (7)
where $U(\mathbf{x}_{t+1})=\max\{u(1,\mathbf{x}_{t+1}),u(0,\mathbf{x}_{t+1})\}$. Because $(S(t),V(t))$ with $V(t)=(n+t)S(t)-\sum_{l=1}^{n+t}(l-1)X_{l}$ is a sufficient statistic for the model, and the utility depends on $\mathbf{X}_{t}$ only through $S(t)$, the quantity $u(0,\mathbf{x}_{t})$ depends on $\mathbf{x}_{t}$ only through $(S(t),V(t))$. The decision tree reduces from exponential in $T$ to a three-dimensional table indexed by $(t,s,v)$, with backward recursion:
$u_{t}(0,s,v)=\sum_{h=0}^{\infty}U_{t+1}(s+h,v+s+h)\\ \times P(S(t+1)=s+h,V(t+1)=v+s+h\mid S(t)=s,V(t)=v).$ (8)
The transition probabilities are estimated by Monte Carlo simulation from the posterior predictive distribution. In practice, the table can be further reduced to two dimensions $(t,s)$ when the accumulation curve is near its plateau and $V(t)$ is approximately determined by $S(t)$.
Carlin et al. [1998] applied this sufficient-statistic reduction to clinical trials directly, developing sequential decision-theoretic methods for treatment allocation. Their approach demonstrated that backward induction—the “theoretically sound and generally accepted procedure for Bayesian sequential analysis” that is “hardly ever implemented due to its great complexities” [Berger, 1985]—becomes feasible when the problem structure is exploited.
3.4 The Normal model and gridding: a worked example
We develop the backward induction in full for the clinical trial problem of Brockwell and Kadane [2003], following Carlin et al. [1998]. This example makes the sufficient-statistic reduction concrete and illustrates the gridding algorithm that makes optimal sequential design computationally tractable.
Setup.
Patients arrive sequentially, and the outcome $Z_{k}$ for patient $k$ given treatment effect $\theta$ satisfies
$Z_{k}\mid\theta\sim N(\theta,\sigma^{2}),\qquad\theta\sim N(0,\sigma_{0}^{2}),$ (9)
with $\sigma^{2}$ known. The decision at each stage is either to stop and choose a terminal action $d\in\{d_{0},d_{1}\}$ (declare for control or treatment) or to continue sampling at cost $c$ per patient. The loss for terminal decision $d_{1}$ (favor treatment) when the true effect is $\theta$ is $R(\theta,d_{1})$, and similarly for $d_{0}$; a natural choice is $R(\theta,d_{1})=-\theta$ (reward equals effect size if treatment is chosen) and $R(\theta,d_{0})=0$.
Sufficient statistics.
After $n$ observations, the posterior is $\theta\mid z_{1},\ldots,z_{n}\sim N(S_{n},\psi_{n}^{2})$, where
$S_{n}=E[\theta\mid z_{1},\ldots,z_{n}]=\psi_{n}^{2}\left(\frac{S_{n-1}}{\psi_{n-1}^{2}}+\frac{z_{n}}{\sigma^{2}}\right),\qquad\psi_{n}^{2}=\left(\frac{1}{\psi_{n-1}^{2}}+\frac{1}{\sigma^{2}}\right)^{-1}=\frac{\sigma^{2}\psi_{n-1}^{2}}{\sigma^{2}+\psi_{n-1}^{2}}.$ (10)
Since $\psi_{n}^{2}$ is deterministic (it depends only on $n$, $\sigma^{2}$, and $\sigma_{0}^{2}$), the state of the trial after $n$ patients is fully described by the scalar $S_{n}$. Decision rules, expected losses, and continuation values all depend on data only through $S_{n}$.
Predictive transitions.
The next observation has predictive distribution
$Z_{n+1}\mid S_{n}\sim N(S_{n},\,\sigma^{2}+\psi_{n}^{2}),$ (11)
and since $S_{n+1}$ is a linear function of $Z_{n+1}$, the updated posterior mean is also Gaussian:
$S_{n+1}\mid S_{n}\sim N\!\left(S_{n},\;\frac{\psi_{n}^{4}}{\sigma^{2}+\psi_{n}^{2}}\right).$ (12)
Equations (10)–(12) show that the entire sequential problem can be written as a Markov chain on the one-dimensional sufficient statistic $S_{n}$, with known Gaussian transitions.
Bellman equation.
Let $h_{n}(s)$ denote the expected loss of the optimal terminal decision at stage $n$ when $S_{n}=s$:
$h_{n}(s)=\min_{d\in\{d_{0},d_{1}\}}E[R(\theta,d)\mid S_{n}=s].$ (13)
For the loss $R(\theta,d_{1})=-\theta$, $R(\theta,d_{0})=0$, this gives $h_{n}(s)=\min(0,-s)$: stop and choose treatment if $s>0$, control if $s\leq 0$. Define $V_{n}(s)$ as the optimal expected loss-to-go from state $s$ at stage $n$, with horizon $T$. The Bellman recursion is
$V_{n}(s)=\min\!\Big\{h_{n}(s),\;\;c+\int_{-\infty}^{\infty}V_{n+1}(s^{\prime})\,p(s^{\prime}\mid S_{n}=s)\,ds^{\prime}\Big\},$ (14)
with terminal condition $V_{T}(s)=h_{T}(s)$. The first argument is the loss from stopping now; the second is the cost of one more patient plus the expected future optimal loss, with the transition density $p(s^{\prime}\mid S_{n}=s)$ given by (12). The optimal policy stops at stage $n$ whenever $h_{n}(s)$ is smaller than the continuation value.
Gridding algorithm.
Brockwell and Kadane [2003] evaluate (14) on a fine grid $\mathcal{G}=\{s_{1},\ldots,s_{G}\}$ over the range of $S_{n}$. In their clinical trial example, $\mathcal{G}$ spans [6,6][-6,6] with G=12, [^0] G=12{,}000 points. The algorithm proceeds backward from $T$:
  1. Step 1.
    At $n=T$, set $V_{T}(s_{j})=h_{T}(s_{j})$ for all $s_{j}\in\mathcal{G}$.
  2. Step 2.
    For $n=T-1,T-2,\ldots,0$: compute the continuation value
    $Q_{n}(s_{j})=c+\sum_{s_{i}\in\mathcal{G}}V_{n+1}(s_{i})\,p(s_{i}\mid S_{n}=s_{j})\,\Delta s$
    by numerical quadrature over the Gaussian transition (12). Set $V_{n}(s_{j})=\min\{h_{n}(s_{j}),Q_{n}(s_{j})\}$.
Because $\psi_{n}^{2}$ is deterministic, the transition density width changes with $n$ but is the same for all $s_{j}$ at a given $n$, so the quadrature weights can be precomputed once per stage.
Triangular stopping regions.
The optimal policy partitions the $(n,S_{n})$ plane into three regions: stop and choose treatment ($S_{n}$ large), stop and choose control ($S_{n}$ small), and continue (intermediate $S_{n}$). These regions form a triangle that narrows with $n$: early in the trial, the continuation region is wide (much can still be learned); as $n$ approaches $T$, the boundaries converge and the trial is forced to a decision. Figure 4 shows the result for the Normal model with $\sigma^{2}=\sigma_{0}^{2}=1$ and sampling cost $c=0.02$, together with six simulated trial paths that terminate when they cross a boundary. This is the Bayesian analog of the triangular continuation region in frequentist group sequential designs [Pocock, 1977], but here the boundaries are derived from the loss function and the prior rather than from alpha-spending constraints.
Figure 4: Optimal stopping regions from backward induction on the Normal model (9) with $\sigma^{2}=\sigma_{0}^{2}=1$ and sampling cost $c=0.02$ per patient. The upper region (green) indicates stopping in favor of treatment; the lower region (red) indicates stopping in favor of control; the middle region indicates continuing. Colored curves show six simulated trial paths with different true $\theta$ values, each terminating when it crosses a boundary.
Extensions.
When $\sigma^{2}$ is unknown, $\psi_{n}^{2}$ is no longer deterministic and must be tracked as a second sufficient statistic; adding $S_{n}^{(2)}=E[\theta^{2}\mid\text{data}]$ restores the Markov property, and the grid becomes two-dimensional. For multi-arm trials, each arm has its own posterior mean, and the grid dimension equals the number of arms. The computational cost grows polynomially in the grid size and linearly in the horizon, making problems with 2–3 arms and horizons of several hundred patients tractable on modern hardware.
The connection to $e$-values is worth noting. Grünwald et al. [2024] develop hypothesis testing based on $e$-values—nonnegative random variables $E$ satisfying $E_{P}[E]\leq 1$ for all $P\in\mathcal{H}_{0}$—that can be multiplied across sequential analyses while preserving type-I error guarantees. Growth-rate-optimal $e$-variables take the form of Bayes factors with special priors, connecting Bayesian sequential design to modern safe testing theory.
4 Decision-Theoretic Trial Design
A clinical trial involves a sequence of decisions: which dose to administer, which patients to enroll, when to stop, and what to conclude. The frequentist framework treats some of these as design parameters fixed before the trial and others as inference conducted after it. The Bayesian decision-theoretic framework treats them uniformly as decisions under uncertainty, optimized with respect to a utility function.
4.1 Utility-based dose-finding: the EffTox design
Thall and Russell [1998] and Thall and Cook [2004] developed designs that jointly model efficacy and toxicity, selecting doses to maximize a clinical utility function rather than merely identifying the maximum tolerated dose.
Let $\pi_{E}(x)$ and $\pi_{T}(x)$ denote the probabilities of efficacy and toxicity at dose $x$. The marginal models use logistic regression:
$\text{logit}[\pi_{E}(x)]$ $=\mu_{E}+\beta_{E,1}x+\beta_{E,2}x^{2},$ (15)
$\text{logit}[\pi_{T}(x)]$ $=\mu_{T}+\beta_{T}x.$ (16)
The quadratic term in (15) allows for a non-monotone dose-efficacy relationship (efficacy may plateau or decline at high doses), while toxicity is assumed monotone increasing.
A utility function $U(\pi_{E},\pi_{T})$ is elicited from clinicians by asking them to specify trade-off contours: combinations of $(\pi_{E},\pi_{T})$ that are equally desirable. For example, a clinician might judge that 60% efficacy with 0% toxicity is equivalent to 100% efficacy with 30% toxicity. These elicited points define indifference curves in the efficacy-toxicity plane.
At each stage, the design:
  1. 1.
    Updates the posterior distribution of all model parameters given accumulated data.
  2. 2.
    Identifies admissible doses: those with posterior probability exceeding a threshold that efficacy exceeds a minimum $A_{E}$ and toxicity is below a maximum $A_{T}$.
  3. 3.
    Among admissible doses, selects the one maximizing posterior expected utility
    $E[U(\pi_{E}(x),\pi_{T}(x))\mid\text{data}]$.
This framework generalizes naturally. Thall [2024] extends it to precision medicine, where utility functions and dose-finding rules are tailored to prognostic subgroups defined by biomarkers.
4.2 Adaptive enrichment
When a treatment benefits only a biomarker-defined subgroup, an enrichment design modifies enrollment criteria based on interim data. Starting with broad enrollment, the trial evaluates treatment effects within pre-specified subgroups. If evidence accumulates that the treatment is ineffective in certain subgroups, enrollment is restricted to the responsive population.
The decision at each interim analysis—whether to continue broad enrollment, restrict to a subgroup, or stop entirely—is a multi-armed bandit problem where the “arms” are enrollment strategies rather than treatments. The posterior predictive probability of eventual success in each subgroup guides the decision.
5 Case Studies
5.1 ECMO: the role of priors with minimal data
The 1985 trial of extracorporeal membrane oxygenation (ECMO) for newborns with persistent pulmonary hypertension [Bartlett et al., 1985] used a randomized play-the-winner design [Wei and Durham, 1978]: adaptive randomization that favors the more successful treatment. The first patient was randomized to ECMO and survived; the second to conventional medical therapy (CMT) and died; the remaining ten were all randomized to ECMO and all survived. Final tally: 11/11 ECMO survivors, 0/1 CMT survivors.
Ware [1989] discussed the statistical challenges; Kass and Greenhouse [1989] provided the Bayesian perspective. With a uniform prior $\text{Beta}(1,1)$ on each treatment’s survival probability, the posteriors are:
$\theta_{\text{ECMO}}\mid\text{data}$ $\sim\text{Beta}(12,1),\quad E[\theta_{\text{ECMO}}\mid\text{data}]=12/13\approx 0.923,$ (17)
$\theta_{\text{CMT}}\mid\text{data}$ $\sim\text{Beta}(1,2),\quad E[\theta_{\text{CMT}}\mid\text{data}]=1/3\approx 0.333.$ (18)
The posterior probability $P(\theta_{\text{ECMO}}>\theta_{\text{CMT}}\mid\text{data})$ can be computed exactly via Thompson’s formula (3) and exceeds 0.99.
The critical issue is sensitivity to the prior. With only one control patient, the posterior for $\theta_{\text{CMT}}$ is dominated by the prior. Historical data suggested approximately 20% survival with CMT from large observational series. Encoding this as $\text{Beta}(4,16)$—equivalent to 20 prior patients with 4 survivors—yields $\theta_{\text{CMT}}\mid\text{data}\sim\text{Beta}(4,17)$ with posterior mean 0.19, and $P(\theta_{\text{ECMO}}>\theta_{\text{CMT}}\mid\text{data})$ remains above 0.99. The conclusion is robust because the ECMO evidence (11/11) is overwhelming; the prior matters only for the magnitude of the estimated difference, not the direction.
Berry [1989] emphasized that the Bayesian framework naturally handles the controversial stopping rule: under the likelihood principle, inference depends on the data actually observed, not on the stopping rule that generated them. A p-value for the same data depends on whether the investigator planned to stop after 12 patients, planned to stop after one CMT death, or had no fixed plan—the Bayesian posterior does not.
The ECMO trial was reanalyzed in the modern EOLIA trial for adult ARDS [Goligher et al., 2018]. The frequentist p-value was 0.09 (not significant at the conventional 0.05 level), but Bayesian analyses with various priors showed 88–99% posterior probability that ECMO reduces mortality. This illustrates how the Bayesian framework provides more nuanced and actionable conclusions than a binary significant/not-significant determination.
5.2 CALGB 49907: predictive probability monitoring
The Cancer and Leukemia Group B trial 49907 tested whether capecitabine could replace standard chemotherapy in women with early-stage breast cancer aged 65 or older [Muss et al., 2009]. The planned sample size ranged from 600 to 1800 patients, with Bayesian predictive probability monitoring for futility and noninferiority at pre-specified interim analyses.
Capecitabine would be declared noninferior if the hazard ratio for disease recurrence or death exceeded 0.8046 (corresponding to a 5-year relapse-free survival of 53% vs. 60% for standard therapy). At each interim analysis, the predictive probability was computed: given current data, what is the probability that future follow-up will yield a conclusive result?
At the first interim analysis after 600 patients, the hazard ratio was 0.53 (capecitabine markedly inferior). The Bayesian probability that this ratio was less than 0.8046 was 96%, exceeding the futility threshold of 80%. Enrollment was discontinued. After an additional year of follow-up, the hazard ratio was 2.09 (95% CI: 1.38–3.17; $P<0.001$), confirming that standard chemotherapy is superior.
The trial achieved its scientific goal with 633 patients instead of the 1800 maximum—a savings of 65% in sample size and years of accrual time. This was the first fully Bayesian randomized trial published in the New England Journal of Medicine [Muss et al., 2009]. The design used noninformative priors for treatment effects; the Bayesian contribution was entirely in the decision framework (predictive probabilities for stopping), not in the priors.
5.3 Platform trials and the Bayesian Time Machine
Platform trials test multiple treatments within a single protocol, with staggered entry and exit of experimental arms sharing a common control. A central question is whether analyses for a given treatment arm should use only concurrent controls or also non-concurrent controls enrolled earlier in the platform.
Saville et al. [2022] proposed the Bayesian Time Machine: a hierarchical model that smooths control-arm response rates across time, allowing analyses to use all control data while adjusting for temporal drift. The model includes time effects as random effects with a smoothing prior that penalizes abrupt changes, so that closer time intervals are modeled with similar response rates.
In simulations, the Bayesian Time Machine provides estimates with smaller mean squared error than either concurrent-only or pooled-controls analyses. The smoothing prior acts as a compromise: when temporal drift is absent, it borrows efficiently across all time periods; when drift is present, it limits borrowing to adjacent periods. The I-SPY 2 platform trial for neoadjuvant breast cancer employed this approach across 23 experimental therapies, of which nine graduated to Phase III based on exceeding an 85% Bayesian predictive probability of confirmatory trial success [Berry, 2025].
6 The Regulatory Landscape
The FDA’s January 2026 draft guidance on Bayesian methodology [Food and Drug Administration, 2026] represents a significant shift from philosophical debate to practical specification. The guidance requires:
  • Pre-specification of Bayesian success criteria and decision thresholds.
  • Evaluation of operating characteristics through simulation, including type I error under the null hypothesis.
  • Justification of prior distributions, with quantification of their influence (e.g., effective sample size).
  • Sensitivity analyses examining robustness to prior specification and prior-data conflict.
  • Documentation sufficient for regulatory reproducibility.
The guidance identifies specific settings where Bayesian methods have supported successful submissions: borrowing from previous trials (e.g., REBYOTA for recurrent C. difficile), augmenting concurrent controls with external data (e.g., GBM AGILE, Precision Promise), pediatric extrapolation (e.g., empagliflozin, linagliptin for pediatric type 2 diabetes), and dose-finding trials in oncology.
A common misconception, noted even in the draft ICH E20 guideline on adaptive trials, equates Bayesian methods primarily with informative priors for borrowing. The majority of successful Bayesian adaptive trials use minimally informative priors overwhelmed by accumulating data. The benefits come from the coherent inferential framework, the interpretability of posterior probability statements, and predictive probabilities—which have no frequentist analog.
7 Discussion
The three themes of this paper—prior specification, sequential design, and decision-theoretic optimization—are not independent ideas that happen to share a Bayesian label. They are consequences of a single principle: probability is the language of uncertainty, and decisions under uncertainty should maximize expected utility.
When this principle is applied to clinical trials, priors are not optional nuisance parameters but expressions of genuine knowledge. The fact that treatment effects are small is prior information as real as the trial data. Ignoring it (via Jeffreys’ or uniform priors) is not objectivity—it is a specific, and poor, modeling choice. The Pólya-Gamma augmentation of Polson et al. [2013] removes the historical computational barrier to exact Bayesian inference for the logistic models most commonly used in trial analysis.
Sequential allocation follows because the posterior updates after every patient, and each update should inform the next decision. Thompson’s (1933) allocation rule is the simplest expression of this idea; the backward induction framework of Christen and Nakamura [2003] and Carlin et al. [1998] provides the optimal one, made tractable by the sufficient-statistic reduction. The predictive probability approximation of Marion et al. [2025] makes real-time monitoring computationally practical.
Decision-theoretic design follows because clinical trials involve multiple objectives—efficacy, toxicity, information gain, cost—that must be traded off explicitly. The utility-based framework of Thall and Cook [2004] demonstrates that this tradeoff can be formalized and optimized, producing designs that are simultaneously more ethical (fewer patients on inferior treatments) and more informative (better identification of optimal doses and subgroups).
The regulatory environment has evolved to accommodate these methods. The 2026 FDA guidance [Food and Drug Administration, 2026] provides a practical roadmap. The I-SPY 2 platform trial [Berry, 2025], the CALGB 49907 trial [Muss et al., 2009], and numerous adaptive designs in oncology and rare diseases demonstrate that Bayesian methods deliver on their theoretical promise in practice.